A spherical balloon is inflated so that its volume is increasing at the rate of 3 ft 3/min. How fast is the radius of the balloon increasing when the radius is 1 ft?
Answer #1


Screen Shot 2021-10-06 at 2.32.51 PM.png

answered by: Kryzha
Answer #2

SOLUTION :


Volume of ballon, V = 4/3 π r^3  


Differentiating w.r.t. t :


=> dV/dt = 4/3 π * 3 r^2 dr/dt 


dV/dt at r = 1 ft. Is 3 ft^3 / min 


=> 3 = 4/3 π  * 3 * (1)^2  dr/dt 

=> 3 = 4 π dr/dt 

=> dr/dt = 3/(4 π)  

=> dr/dt = 0.2387 ft / min 


So, radius is increasing at the rate of 0.2387 ft / min when radius is 1 ft . (ANSWER).


answered by: Tulsiram Garg